Week of 160411

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Maarten (ALICE), Stefan (LHCb), Marcelo (LHCb), Cedric (ATLAS), Julia, Andrea (FTS), Fernando (Compute)
  • remote: Michael (BNL), Dimitri (KIT), Dmytro (NDGF), Rolf (IN2P3), John (RAL), Antonio (CNAF), Jose (PIC), Sang (KISTI)

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • MC15c (digi+reco reconstruction) ongoing
      • HeavyIon reprocessing started last Friday. Using now ~20k slots. Monitoring memory consumption
    • Problems with with some central services (pilot factories, rucio-ui), now fixed.
    • In the process of closing the consistency check tickets opened last year to ask the sites to provide storage dumps.
    • Identified a list of 17 corrupted RAW files on Castor (GGUS:119750). Problem linked to faulty router corrupting the traffic on June 11th last year : OTG:0022230. Insidious corruption since the adler32 of the corrupted files are the same as the good file ! But md5 do differ !!

  • ALICE -
    • High activity on average.
      • New record 102190 running jobs reached Apr 10.
      • More than 100k jobs for many hours since Sun morning.

  • LHCb reports (raw view) -
    • Site Issues
      • T0: Problem with files accessing EOS Disk, solved during the week-end

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF:
    • Availability Reliability computations requested for Alice and Cms, due to error putting CEs in downtime last week.
    • Atlas low IO queues created at our extension site (ReCas).
  • FNAL:
  • GridPP:
  • IN2P3: internal network outage on Sunday, April 10th, from 12:26 to 12:40 because of an unexpected reboot of the central router, cause still unknown. The batch system was impacted, xrootd for Alice wasn’t working for nearly the rest of the day, dCache (staging part) was out of service for one hour. Other services either continued to work after a short interruption or did a failover to other sites.
  • JINR:
  • KISTI: NTR
  • KIT: NTR
  • NDGF:
    • one subsite has power and cooling maintenance Wed-Thu; that means reduced computing capacity Wed-Thu and reduced storage capacity Thu.
    • a short storage outage between 10-11 UTC on Thu to update/reboot the headnode.
  • NL-T1:
  • NRC-KI:
  • OSG:
  • PIC: NTR
  • RAL: One broken LHCb disk server. Some files are unavailable. The diskserver was repaired and then it broke again, it will take some days to repair.
  • TRIUMF:

  • CERN computing services:
    • Infrastructure Services:
      • NTR
    • Linux & Configuration Support:
      • NTR
    • Monitoring & Messaging:
      • NTR
    • Resource Provisioning Services:
      • NTR
  • CERN storage services: NTR
  • CERN databases:
  • GGUS:
  • Monitoring:
    • Draft reports for the March availability sent around
  • MW Officer:
  • Security:

AOB:

Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2016-04-11 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback